Almighty1:

Irene has a credit that uses the previous balance method. The opening balance of one of her 30-day billing cycles was $2510, but that was her balance for only the first 5 days of the billing cycle, because she then paid off her entire balance and didn't make any new purchases. If her credit card's APR is 11%, which expression could be used to calculate the amount Irene was charged in interest for the billing cycle? \(A. {(\dfrac{0.11}{365}*30)}{($2510)}\) \(B.{(\dfrac{0.11}{365}*30)}{($0)}\) \(C.{(\dfrac{0.11}{365}*30)}\cdot{(\dfrac{(5*$2510+25*0)}{30})}\) \(D.{(\dfrac{0.11}{365}*30})\cdot{(\dfrac{(5*0+25*2510)}{30})}\)

dude:

Okay In order to calculate interest use the following equation \((\dfrac{rate(\%)}{365}\cdot billing~cycle~days)\cdot(Starting~amount)\)